OpenStudy (anonymous):

whats 46 over 1012 im simplest form

OpenStudy (anonymous):

are you kidding me?????????

OpenStudy (anonymous):

how old are you? because if youre like 7, then i apologize.

OpenStudy (anonymous):

im 12

OpenStudy (anonymous):

well. i guess i was assuming youre in the age group of almost everyone else.

OpenStudy (anonymous):

sorry anyways. so, 42/1012 what is their most common factor?

OpenStudy (anonymous):

21 / 506

OpenStudy (anonymous):

divide top and bottom by 46

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@ehuman i dont think so. 46 cant be facotres by 12.

OpenStudy (anonymous):

*factored

OpenStudy (anonymous):

46/46=1 1012/46=22

OpenStudy (anonymous):

i think my answer is correct.

OpenStudy (anonymous):

46/1012 = 1/22 = .0454545455 check it with a calculator.

OpenStudy (anonymous):

oooh. i see. haha. it's 42/1012, not 46. sorry, my bad!

OpenStudy (anonymous):